LASCO's doors are presently closed ahead of maneuvers on 2004/03/30 consisting of momentum management, station keeping, and 180 deg roll. Doors are expected to reopen at ~20:00 UT on 2004/03/31. SOHO is now in a telemetry keyhole (2004/03/19 - 2004/04/04). LASCO data will consist of typically two (occasionally three) continuous periods of real-time data per day. Beginning at 00:00 UT on March 18, 2004 EIT will be in CCD bakeout mode; No EIT images until April 5, 2004.
S West Gusty outflow along the streamer all along the day. 04:40 UT E Limb Faint and narrow spray front. It fades in C2. 05:40 UT S West At PA coincident with that of the small ragged fronts on previous day (PA ~240), faint ragged loop front. 07:00 - 09:20 UT Data Gap. 09:20 UT S East Faint ragged loop front that fades in C2. 11:00 - 15:00 UT Data Gap. 15:20 UT S West Ragged loop front already in progress after data gap. 20:20 UT S West Narrow ragged loop front. 21:40 - 24:00 UT Data Gap.
S West Gusty outflow along the streamer all along the day. 00:40 UT S East Bright, small and elongated feature, followed at 05:20 by a similar one. Extremely faint ragged front develops by the time of the latter, fading in C2. 00:40 UT N East Faint ragged front along streamer already in progress. Barely visible in C3. 07:20 - 14:00 UT Data Gap. 14:20 UT N East Faint loop front already in progress after data gap. Another faint loop front starts developing. 17:40 UT S East Bright ragged loop front immediately preceded by a small and elongated feature similar to that one reported at 00:40 UT. Infalling material by the time, slightly to S.
00:20 UT E Limb Slow evolution of ragged front is seen to develop after following data gap as a big ragged loop front with structured interior. Very well seen in C3. 02:40 - 09:20 UT Data Gap. 10:00 UT S East Loop front develops behind previous event. The resulting complex event (streamer blowout) extends toward southern latitudes as it develops. 12:40 - 15:20 UT Data Gap. 15:20 UT S West Couple of ragged loop fronts already in progress after data gap.
00:20 UT N East Narrow ragged loop front followed by a couple more at ~ 07:40 (already in progress after data gap) and 10:00 UT. 00:40 UT S West Bright loop front, strong cavity, and twisted inner core. 01:00 UT S East Bright spray-like front. 03:20 - 07:40 UT Temporary DSN Data Gap. 08:00 UT S East Spray-like front already in progress after data gap. Also couple of ragged fronts roughly above E Limb already in progress. 11:00 UT LASCO's doors close.